OpenStudy (anonymous):

integrate 1/((x^5)*sqrt(9*x^2-1))

OpenStudy (anonymous):

I've gotten pretty far I think, set x= 1/3 sec(U)

OpenStudy (anonymous):

then I got to the point where I had \[81 \int\limits_{}^{} (\cos(x))^4 dx\]

OpenStudy (anonymous):

the I apply the reduction formula twice and sub back in values for x to get (81/4)*(1/(3*x))^(3)*((sqrt(9 x^2-1))/(3*x))+243/(8)*(1/(3*x))*((sqrt(9 x^2-1))/(3*x))+243

OpenStudy (anonymous):